OpenStudy (anonymous):

I can't find critical points for y=x(9-x^2)^1/2

OpenStudy (anonymous):

First take the derivative dy/dx=\[(9-2x ^{2})/\sqrt{9-x ^{2}}\]

OpenStudy (anonymous):

next we must find where dy/dx = 0 or where it is undefined

OpenStudy (anonymous):

dy/dx = 0 for \[x = \pm \sqrt{9/2}\]

OpenStudy (anonymous):

dy/dx is undefined for \[x = \pm3\], when the denominator equals 0

OpenStudy (anonymous):

Are you looking for a specific type of critical point?

OpenStudy (anonymous):

\[y = x \sqrt{9-x^2}\] the derivative of it is: \[y'=\sqrt{9-x^2} -x^2/\sqrt{9-x^2}\] then y'=0 finely x =2.45

OpenStudy (anonymous):

that is not the only answer.

OpenStudy (anonymous):

and the derivative is not quite correct.

OpenStudy (anonymous):

The critical points should be \[x = \pm \sqrt{9/2}\]

OpenStudy (anonymous):

oops, sorry corec, the derivative is fine